OpenStudy (anonymous):

In a certain store the regular price of a refrigerator is $600 how much money is saved by buying this refrigerator at 20 percent off the regular price than buying it on sale at 10 percent off the regular price with an additonal discount of 10 percent off the sale price

OpenStudy (anonymous):

A.$6 B.$12 C.$24 D. $54 E.$60

OpenStudy (anonymous):

please show your work please @satellite73

OpenStudy (danjs):

Calculate both scenarios and compare the difference between the two

OpenStudy (danjs):

20% off of $600, is the same as paying 80% of the price To convert a percent to a decimal, divide by 100. 80% of 600 is (0.80)*600 = $480 paid in the first scenario

OpenStudy (danjs):

Now calculate 10% off, then 10% off that discounted price...

OpenStudy (danjs):

10% off 600*0.9 = $540 Now 10% off of that price. 0.90*540 = $486 paid in the second scneario what is the difference